OpenStudy (anonymous):

Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C) 1/2x + 8 = 10 D) 1/2(2x - 6) = -6

OpenStudy (freckles):

you could plug in -3 for x into equation equation and see if you get the same thing on both sides

OpenStudy (freckles):

for example 2x-7=-1 replace x with -3 2(-3)-7 -6-7 -13 but -13 is not -1 so x=-3 is not a solution to 2x-7=-1 you try the others

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

np